OpenStudy (calebb8133):

Gary used candle molds, as shown, to make candles that were perfect cylinders and spheres: A cylindrical mold is shown, the radius of the top circular section of the cylinder is labeled 2 inches and the height of the cylinder is labeled as 4 inches. On the right side of this mold is a spherical mold. The radius of this spherical mold is labeled as 2 inches. What is the approximate difference in the amount of wax needed to make a candle from each of these molds? Use π = 3.14. 16.75 cubic inches 20.93 cubic inches 24.25 cubic inches 33.49 cubic inches

OpenStudy (calebb8133):

@Will.H

OpenStudy (calebb8133):

@horse_rider_133

OpenStudy (calebb8133):

help i have to get this right!!!!!!

OpenStudy (calebb8133):

i think D

OpenStudy (horse_rider_133):

i dont know it sowwy calebb

OpenStudy (calebb8133):

:(

OpenStudy (calebb8133):

do u know someone who might?

OpenStudy (horse_rider_133):

@tsh

OpenStudy (calebb8133):

@Will.H

OpenStudy (calebb8133):

@triciaal

OpenStudy (calebb8133):

########

OpenStudy (calebb8133):

I FORGOTTEN THIS

OpenStudy (horse_rider_133):

@kj

OpenStudy (calebb8133):

@KJSaif

OpenStudy (calebb8133):

:)

OpenStudy (horse_rider_133):

@KJSaif

OpenStudy (calebb8133):

help plz

KjSaif (kjsaif):

@Will.H

OpenStudy (calebb8133):

i think i is D

OpenStudy (horse_rider_133):

@codes

OpenStudy (horse_rider_133):

@@Lxdrpr

OpenStudy (calebb8133):

thank i @Will.H

OpenStudy (calebb8133):

*U

OpenStudy (horse_rider_133):

he helped u

OpenStudy (will.h):

Alright. Apparently we will have to find the volumes and then find the difference. I'll bring my calculator hold up

OpenStudy (calebb8133):

np

OpenStudy (calebb8133):

i realy want to get an A because then i will be able to get this http://www.horizonhobby.com/products/TRA580641T1

OpenStudy (calebb8133):

*really

OpenStudy (calebb8133):

now i think it is D but i just want to be 100% sure!!!

OpenStudy (calebb8133):

@Will.H where u go?

OpenStudy (will.h):

A cylinder volume is V = pi * r^2 * h We have r as 2 H as 4 V = 50.27 Now find volume of sphere V = (4/3) * pi * r^3 V = (4/3) * pi * 2^3 V = 33.51 Difference is 50.27 - 33.51 = 16.76

OpenStudy (calebb8133):

ok thx :)